OpenStudy (clamin):

ONE MEDAL!! What is F1 and F2 in the punnett square? How to do a RATIOS of the offspring in the punnette square?.. PLEASE HELP!!

OpenStudy (frostbite):

F standing for latin word filial means son or daughter of a specific person or thing. The prenatal generation (P-generation) can be started to set into a punnett square. Assume we got a man with the alleles AA and a woman with aa, we get a punnett square that lo From this square, we assume that A is a dominant, gain a function, allele. So in case of a homozygote dominant subject is crossed with a homozygote recessive subject, the ratio of the offsprings are aA: 4/4 = 100% Lets take any of one the offsprings from the F1 generation and let it cross with another one from the F1 generation, we get a new square  We can see the ratios are: AA: 1/4 = 25% aA: 2/4 = 1/2 = 50% aa: 1/4 = 25%